Fermi National Laboratory
 

Accelerator Update for January 15, 2007 - January 17, 2007

 

Monday January 15

The day shift began with Operations monitoring store 5177 (57.38E30) and a Recycler stash (226.0E10), with the Antiproton Source (Pbar) stacking (18.4E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 8:41 AM, one of the Meson M01 LCW systems tripped off. Operators tried to restart the system, but it tripped off again. Operations contacted experts.

The Neutron Therapy Facility (NTF) reported at 10:53 AM that they were done with the beam for the day.

At 1:18 PM, the Linac Klystron Debuncher tripped for the first of two times in 50 minutes on a cavity LCW flow indication. Operators investigated both times and found that the cavity water pump had tripped off. It reset both times.

Operations established beam to the Switchyard (SY) absorber at 2 PM.

The Recycler completed its third and last stash for the day shift at 3:43 PM.

At 4:50 PM, MI-10 suffered a vacuum burst. The vacuum recovered in 10 minutes.

Operations terminated store 5177 at 5:05 PM.

Operations prepared the CDF collision hall for an access at 5:25 PM. They came out of access about 45 minutes later.

At 6:06 PM, a Linac Klystron RF station (KRF6) tripped off due to a NIM module problem. It wouldn’t reset. Operators contacted an expert who suggested changing the waveguide pressure module. This cleared the fault.

At 7:01 PM, the Tevatron (TeV) suffered a sector E2 quench. Cryo system technicians had worked on a wet engine and they think that there may have been some warm helium gas in the engine when it was turned on and that caused the quench.

At 7:23 PM, the Linac buncher and vernier tripped off for the first of three times in 30 minutes on a cavity vacuum fault. It reset each time.

At 8:45 PM, the stacking rate dropped due to a Pbar Debuncher RF station (DRF1-2) trip as well as some other equipment trips. Operators eventually got everything on and the stack rate back to normal.

Operations put the TeV into shot setup at 9:55 PM.

Operations established store 5179 at 11:33 PM with an initial luminosity of 236.91E30.

Tuesday January 16

The midnight shift began with Operations monitoring store 5179 (213.08E10) and the stash (15.8E10), with Pbar stacking (76.7E10), and with NuMI taking beam.

At 12:15 AM, a Booster vacuum pump tripped off and caused two RF stations (BRF7 & 8) to trip. Everything reset and beam returned a minute later.

The Recycler made its last of three stashes for the midnight shift at 6:33 AM.

At 6:42 AM, the Booster East Anode Power Supply tripped off on a current input over-limit. Operators had the power supply reset and on 10 minutes later.

At 7:42 AM, store 5179 aborted on a TeV sector A2 volts-to-ground indication. Later, the cause was found to be a blown fuse. There was no quench. Operations contacted experts.

A MI expert began a study at MI-60 at 9:47 AM.

Operations put the TeV through a dry squeeze at 9:54 AM.

TeV experts began a study period at 10:37 AM.

At 12:33 PM, Operations reported that the NuMI horn had tripped off three times on a ground fault indication. Operations contacted an expert. NuMI resumed taking beam about 45 minutes later.

Operations put the TeV into a shot setup at 1:02 PM.

Operations established store 5182 at 2:37 PM with an initial luminosity of 195.87E30.

Operations reported at 4 PM that the Recycler had stashed twice during the day shift.

At 10:11 PM, the NuMI beam permit tripped off due to a beam position problem. It reset.

At 10:16 PM, NuMI tripped off for the first of four times in seven minutes on a horn stripline ground fault. It reset each time.

Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.

Wednesday January 17

The midnight shift began with Operations monitoring store 5182 (66.77E30) and the stash (173.1E10), with Pbar stacking (13.3E10), and with NuMI taking beam.

At 1:39 AM, BRF6 tripped off for the first of two times in 90 minutes. The station reset both times.

At 7:02 AM, NuMI tripped off due to a beam position problem. It reset.

Operations reported at 8 AM that that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/17/2007    email Fermilab